OpenStudy (anonymous):

sin^2 theta = 4-2cos^2theta Solve.

OpenStudy (anonymous):

\[sin^2\theta = 1-\cos^2\theta \] use this as a substitution for the left hand side of this equation, then you will obtain an easier equation.

OpenStudy (anonymous):

yeah I got that far, then what….

OpenStudy (anonymous):

\[cos^2\theta=3 \] \[cos \theta = \sqrt{3}\] you could take the inverse function of that, but we know already that there is no solution for that because it's only defined between values of -1 and 1

OpenStudy (anonymous):

OKAY Thank youuuu

OpenStudy (anonymous):

you're welcome